Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

In a certain factory, the production line which produces...

Expert replies
by BTGmoderatorLU » Thu Jan 04, 2018 8:34 am
In a certain factory, the production line which produces the bags has the probability that a bag selected at random is defective is 0.02. If 6 bags are selected at random, what is the probability that at least one bag is defective?

$$A.\ 0.02^6$$
$$B.\ \left(0.02\right)\left(0.98\right)^6$$
$$C.\ \ 1-\left(0.02\right)^6$$
$$D.\ \ 0.98^6$$
$$E.\ \ 1-0.98^6$$

The OA is E.

I'm really confused with this PS question. Experts, any suggestion about how can I solve it? Thanks in advance.
Join the discussion
Source: — Problem Solving |

by Brent@GMATPrepNow » Thu Jan 04, 2018 8:43 am
LUANDATO wrote:In a certain factory, the production line which produces the bags has the probability that a bag selected at random is defective is 0.02. If 6 bags are selected at random, what is the probability that at least one bag is defective?

$$A.\ 0.02^6$$
$$B.\ \left(0.02\right)\left(0.98\right)^6$$
$$C.\ \ 1-\left(0.02\right)^6$$
$$D.\ \ 0.98^6$$
$$E.\ \ 1-0.98^6$$
We want P(select at least 1 defective bag)
When it comes to probability questions involving "at least," it's best to try using the complement.
That is, P(Event A happening) = 1 - P(Event A not happening)
So, here we get: P(getting at least 1 defective bag) = 1 - P(not getting at least 1 defective bag)
What does it mean to not get at least 1 defective bag? It means getting zero defective bags.
So, we can write: P(getting at least 1 defective bag) = 1 - P(getting zero defective bags)

ASIDE: If P(bag is defective) = 0.02, then P(bag is NOT defective) = 0.98

P(getting zero defective bags)
P(getting zero defective bags) = P(1st bag is good AND 2nd bag is good AND 3rd bag is good AND 4th bag is good AND 5th bag is good AND 6th bag is good)
= P(1st bag is good) x P(2nd bag is good) x P(3rd bag is good) x P(4th bag is good) x P(5th bag is good) x P(6th bag is good)\
= 0.98 x 0.98 x 0.98 x 0.98 x 0.98 x 0.98
= 0.98^6

So, P(select at least 1 defective bag) = 1 - 0.98^6
= E

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by Scott@TargetTestPrep » Mon Sep 02, 2019 6:06 pm
BTGmoderatorLU wrote:In a certain factory, the production line which produces the bags has the probability that a bag selected at random is defective is 0.02. If 6 bags are selected at random, what is the probability that at least one bag is defective?

$$A.\ 0.02^6$$
$$B.\ \left(0.02\right)\left(0.98\right)^6$$
$$C.\ \ 1-\left(0.02\right)^6$$
$$D.\ \ 0.98^6$$
$$E.\ \ 1-0.98^6$$

The OA is E.

I'm really confused with this PS question. Experts, any suggestion about how can I solve it? Thanks in advance.
We are given that the probability that a selected bag is defective is 0.02, so we can determine that the probability of a non-defective bag is 1 - 0.02 = 0.98. We need to determine the probability that when 6 bags are selected at least one bag is defective.

We know the following:

1 = P(at least one defective bag is selected) + P(no defective bags are selected)

Thus,

P(at least one defective bag is selected) = 1 - P(no defective bags are selected)

Therefore, we need to determine P(no defective bags are among the 6 selected bags).

P(no defective bags are among the 6 selected bags) = 0.98^6

Thus, P(at least one defective bag is among the 6 selected bags) = 1 - 0.98^6

Answer: E

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion